My yard is 9 meters long and 8 meters wide. And how many meters of fencing will i need to cover the enitre yard?
Answer by oberobic(2304)

L = 9 m
W = 8 m
.
P = perimeter
P = 2(L+W)
.
P = 2(9+8)
P = 2(17)
P = 34
.
Done.